two angles are supplementary. one angle is 3 degrees less than twice the other. find the measure of the 2 angles
Answer by stanbon(75887)

two angles are supplementary. one angle is 3 degrees less than twice the other. find the measure of the 2 angles
---------------------
Equations:
x + y = 180
x = 2y-3
--------------------
Substitute for "x" and solve for "y":
2y-3 + y = 180
3y = 183
y = 61 degrees
----
x = 2y-3 = 2*61-3 = 119 degrees
==================================
Cheers,
Stan H.
==========